two locomotives approach each other on parallel tracks. Each has a speed of 95 km/h with respect to the ground. If they are initially 8.5 km apart, how long will it be before they reach each other?
Found 2 solutions by oberobic, fcabanski:
Answer by oberobic(2304)

The two locomotives are closing on one another at the sum of their speeds: 95+95 = 190 km/hr.
The start at 8.5 km apart, so they will meet each other pretty soon.
t = 8.5 km/ 190 km/hr
t = 0.0447368421 hr
.
That is a precise but not terribly useful answer.
.
0.0447368421 hr * 60 min/hr = 2.68 minutes

Answer by fcabanski(1391)

Check my profile and contact me for one on one online tutoring.


D=RT.


What distance do they have to travel to reach each other? 8.5 km


What speed are they moving towards each other? 95+95 = 190km/h.


8.5 = 190t ---> divide both sides by 190. 8.5/190=t=.0447hrs, or about 2.682 minutes. Less than 3 minutes.